# NM-FlowGAN: Modeling sRGB Noise with a Hybrid Approach based on Normalizing Flows and Generative Adversarial Networks

### v2

@@ 7 Simultaneous Training
Table 5: Comparison of quantitative performance under different training strategies. PSNR/SSIM values are measured from the SIDD validation dataset.
@@
Our NM-FlowGAN consists of two distinct neural networks: pixel-wise noise modeling networks and spatial correlation networks. These networks are based on Normalizing Flows and GAN, respectively. Our method is capable of training multiple networks simultaneously. However, each networks is optimized independently according to its respective loss function, ensuring that the optimization of one networks does not affect the other networks. In this paper, we indicate this training strategy to simultaneous training. This implementation is possible because each neural networks tends to synthesize a distinct form of noise, and it can improve the training stability and performance of our model in sRGB noise synthesis. Tab. 5 shows the comparison of noise modeling performance for three different training strategies: the simultaneous training strategy, the joint training strategy without using a stop gradient function, and the two-stage training strategy where the pixel-wise and spatial correlation modeling networks are trained sequentially.
In the case of the two-stage training strategy, it shows comparable performances to our proposed simultaneous training strategy. The two-stage training strategy can also be adopted due to our approach of training the neural networks independently. However, in the case of simultaneous training, our use of simple pixel-wise modeling networks leads to rapid convergence of the negative log likelihood (NLL) loss. As a result, the model is trained in a similar manner to the two-step training strategy. Therefore, our proposed simultaneous training strategy and the two-stage training strategy show similar performances. In addition, the training time is extended compared to the other training strategy due to the need for sequential training of each neural network. As a result, we adopt the simultaneous training strategy for efficiency.
8 Detailed Structure of Linear Flow Layers
Fig. 7 illustrates the detailed architecture of the core components of our pixel-wise noise modeling network: the signal-dependent conditional linear flow layer (SDL) and the structure-aware conditional linear flow layer (SAL). These layers are designed to estimate and apply scale and bias values for each pixel using a clean image and conditions, including camera type and ISO setting. The SDL and the SAL first apply a one-hot encoding of the camera type and the ISO setting, and then extract the features by means of an architecture consisting of residual blocks. This method of encoding camera conditions is adopted from sRGBFlow [20]. The extracted features are then channel-wise concatenated with the clean image to serve as input to the feature extraction networks (FEN) of SDL and SAL. Each of these layers has a different architecture for feature extraction, and the outputs from these layers are used as scale and bias parameters for each pixel.
Specifically, the SDL consists mainly of pointwise convolution layers, instance normalization layers, and activation layers. This design aims to estimate log scale and bias parameters based on the clean image intensity of each pixel. On the other hand, the SAL consists mainly of convolution layers with $3\times 3$ filters and activation layers. Its design aims to estimate log scale and bias parameters based on the image structure surrounding each pixel.
8.1 Camera-Specific Noise Modeling
In our main paper, we demonstrate that our NM-FlowGAN effectively models the noise in the SIDD dataset, which consists of images captured by various cameras, by training a single unified model. However, as mentioned, sRGB noise is significantly affected by hardware specifications such as image sensors. In scenarios where the noise modeling data is captured by cameras with widely varying hardware specifications, employing camera-specific training becomes a suitable solution to improve performance.
Table 6: KL divergence of our NM-FlowGAN in camera-specific training. ${}^{\dagger}$ indicates results obtained from camera-specific training. In camera-specific training, each noise modeling networks is trained using only the data that corresponds to each camera type.
Method G4 GP IP N6 S6 ALL
NM-FlowGAN 0.039 0.026 0.017 0.041 0.033 0.028
${}^{\dagger}$ 0.031 0.024 0.010 0.058 0.050 0.032
Tab. 6 demonstrates that our proposed NM-FlowGAN shows promising performance even in camera-specific training. This is due to our hybrid architecture, which exploits the advantages of both Normalizing Flows and GAN, allowing our NM-FlowGAN to be stably trained even with small-sized datasets. Specifically, in our pixel-wise noise modeling networks based on Normalizing Flows, stable training can inherently be achieved even on small-sized datasets. Moreover, our GAN-based spatial correlation modeling networks model spatial correlations from synthesized pixel-wise noise, enabling more consistent modeling compared to other GAN-based noise modeling networks. This consistency leads to more stable training.
9 Latency
Table 7: Average latency of the baselines and our method.
@@ 9 Latency
As shown in Tab. 7, our method not only demonstrates better noise synthesis performance compared to NeCA-W, but also has lower latency. Although our method exhibits a higher latency compared to sRGBFlow, the significant performance improvement brings justifies this increase and makes it tolerable in practice. We measured latency using a batch of 16 images, each with a resolution of $256\times 256$, on a single NVIDIA A6000 GPU.
10 Additional Experimental Results
To further validate the noise modeling performance of our model, we present additional experimental results in this section. First, Fig. 8 shows the standard deviation as a function of clean image intensity for our model (NM-FlowGAN), sRGBFlow [20], NeCA-W [12] and $*{}^{*}$. This figure indicates that our model shows similar behavior to real noise compared to the other baselines. In addition, Fig. 9 shows additional visual results of synthetic noisy and noise images from our noise model and various baselines, and Fig. 10 shows additional visual results from a denoiser trained on synthesized noisy-clean image pairs generated by each noise model. In the figures, it can be observed that our method shows superior performance compared to other baseline methods.
11 Noise Synthesis Pipelines
In our experiments, we present two distinct sets of results for NeCA-W. This distinction arises due to discrepancies between the noise synthesis pipeline described in the method section of the NeCA paper [12] and the noise synthesis pipeline described in the experiments section of the NeCA paper.
In the method section of the NeCA paper, it is mentioned that NeCA-W uses arbitrary noisy images to estimate gain maps. However, in the experiments section of the NeCA paper, they use paired noisy images, each corresponding to a clean image, to eliminate randomness in the experiments. However, this may lead to an overestimation of the noise modeling performance. This is because the real noise image, which is the target of the noise synthesis, can simply be obtained from the difference between the paired clean and noisy images used as inputs to NeCA-W. This means that NeCA-W can consider ground truth data during noise synthesis. Furthermore, paired noisy images may not be achievable in real-world noise synthesis scenarios.
For these reasons, we include both results in our experiments: one with arbitrary noisy images obtained with the same camera settings (e.g. camera type and ISO setting) as described in the method section of the NeCA paper (NeCA-W), and the other with paired noisy images ($*{}^{*}$). To clarify our experimental settings, we provide a comparison of the noise synthesis pipelines for NeCA-W, $*{}^{*}$, sRGBFlow [20], and ours in Fig. 11.
Figure 9: Visualization of synthetic noisy and noise images from our NM-FlowGAN and other baselines under various camera types and ISO settings. This shows that the synthetic noisy and noise images synthesized by our method are visually closer to real noise than those synthesized by other baselines. In addition, in terms of KL divergence measurement against the real noise, our proposed method quantitatively outperforms the other baselines.
Figure 10: Visual comparison of denoising sRGB images in the SIDD validation dataset. Each denoiser is trained on synthesized noisy-clean image pairs generated by each model.
Figure 11: An illustration of noise synthesis pipelines of baselines and ours.

### v3

@@ IV-G Simultaneous Training
TABLE VII: Comparison of quantitative performance under different training strategies. PSNR/SSIM values are measured from the SIDD validation dataset.
@@
In the case of the two-stage training strategy, it shows comparable performances to our proposed simultaneous training strategy. The two-stage training strategy can also be adopted due to our approach of training the neural networks independently. However, in the case of simultaneous training, our use of simple pixel-wise modeling networks leads to rapid convergence of the negative log likelihood (NLL) loss. As a result, the model is trained in a manner similar to the two-stage training strategy. Therefore, our proposed simultaneous training strategy and the two-stage training strategy show similar performances. In addition, the training time is longer compared to the other training strategy because each neural network needs to be trained sequentially. Therefore, we adopt the simultaneous training strategy for efficiency.
IV-H Latency
TABLE VIII: Average latency of the baselines and our method.
@@ IV-H Latency
As shown in Table VIII, our method not only demonstrates better noise synthesis performance compared to NeCA-W, but also has lower latency. Although our method exhibits a higher latency compared to sRGBFlow, the significant performance improvement brings justifies this increase and makes it tolerable in practice. We measured latency using a batch of 16 images, each with a resolution of $256\times 256$, on a single NVIDIA A6000 GPU.
V Conclusion
In this paper, we introduce a novel hybrid method for modeling sRGB noise. We begin by analyzing the characteristics of sRGB noise and exploring the factors that affect its distribution. From this analysis, we present NM-FlowGAN, noise modeling neural networks that combine two generative models: Normalizing Flows and GAN. By leveraging the strengths of both networks, our method combines pixel-wise noise modeling with Normalizing Flows and spatial correlation modeling with GAN. This combination allows our method to effectively capture noise characteristics and model pixel-to-pixel relationships in sRGB noise. In particular, our approach synthesizes noisy images using only clean images and noise-related factors like camera type or ISO settings, thereby enhancing its applicability in scenarios where paired data is unavailable. Our experimental results show that our method outperforms other baseline methods in sRGB noise modeling and image denoising. Furthermore, we demonstrate the effectiveness of our proposed method and its applicability in real-world scenarios through various experiments.